OpenStudy (anonymous):

please quick help with this question ?!| What is the length of the base of a triangle with height 11 cm and area 137.5 cm2?

hero (hero):

\(A_{\triangle} = \dfrac{bh}{2}\) where A = area of the triangle b = base of the triangle h = height of the triangle

hero (hero):

@Autobee7128, you are given the area and height of the triangle and asked to find the base of the triangle, right?

OpenStudy (anonymous):

Yeah i need to find the base

hero (hero):

Okay, so the first thing we should do is isolate b. Do you know how to do that?

OpenStudy (anonymous):

not sure

hero (hero):

To isolate b, Multiply both sides by 2, then divide both sides by h. Upon doing so, you get: \(\dfrac{2A_{\triangle}}{h} = b\)

hero (hero):

Right @Autobee7128? Are you still here?

OpenStudy (anonymous):

yeah so whats the next step?

hero (hero):

The next step is to plug in the given values for the Area and the height, then solve for b. In this case, A = 137.5 h = 11

hero (hero):

Would you like to plug those values in and calculate b?

hero (hero):

Because, you know, after plugging those values in, you'd have \(\dfrac{2(137.5)}{11} = b\) Then you could find the value of b using a calculator. Also, remember, the units for b is cm.

OpenStudy (anonymous):

I was just about to do that. But thanks anyways !! You were very helpful :)

hero (hero):

You're welcome. By the way, what value did you get for b?

OpenStudy (anonymous):

25 !